OpenStudy (anonymous):

determine the series 4/square root of ((K^3)+1) converge or not? thx

OpenStudy (anonymous):

\[\sum_{k=1}^{\infty}4/\sqrt{k ^{3}+1}\]

OpenStudy (anonymous):

it converges to 0 ( because \[ \lim_{k \rightarrow +\infty } (4/\sqrt{k+1}^{3}) = 0\]

OpenStudy (anonymous):

how can it converge to zero? all the terms are positive

OpenStudy (anonymous):

it does converge though

OpenStudy (anonymous):

idea is the degree of the denominator is bigger than 1 ok it is not a polynomial but \[\sqrt{k^3}=k^{\frac{3}{2}}\] and \[\frac{3}{2}>1\] so converges probably use comparison test